# Control of a commercial vehicle by a tetraplegic human using a bimanual brain-computer interface

### v1

@@ DISCUSSION
The neural recordings in this study were collected from one participant enrolled in a BCI clinical study (ClinicalTrials.gov Identifier: NCT01958086). The institutional review boards of the California Institute of Technology, Casa Colina Hospital and Centers for Healthcare, and the University of California, Los Angeles, approved study procedures, including informed consent, implant surgery, and experiment design. BCI-JJ is a right-handed male 50 years old with a C4-C5 level spinal cord injury that occurred approximately three years before his enrollment in the study. BCI-JJ was implanted with two arrays of 96-channel NeuroPort Electrodes, including one array in PPC and the other in MC. Neural signals were acquired, amplified, bandpass filtered (0.3 Hz to 7.5 kHz), and digitized (30 kHz, 16 bits/sample) from electrodes using the neural signal processors called NeuroPort Systems (Blackrock Neurotech, Salt Lake City, UT).
It is important to note that BCI-JJ’s electrodes had been implanted for more than five years at the time of testing. The signal-to-noise ratio (SNR) of the recorded brain signals had decreased to approximately one-third of the value observed in the first year after implantation. Due to this decline, a custom feature extraction algorithm, FENet (see the “BCI decoder system” section), played a crucial role in enabling BCI-JJ to control the vehicle with the proficiency described in this paper.
@@
We recruited 20 motor intact individuals, including 10 males and 10 females with an average age of $50\pm 15$, the same as the age of BCI-JJ at the time of testing. The motor intact control group participated in the simple reaction time task, the braking reaction time task, and the simulated town driving task. Each motor intact participant attended two sessions on their own to complete these three tasks. All participants were informed about the purpose of the study and provided their written informed consent prior to participating. They were compensated at the end of each session based on their task performance.
@@
We used PLSR to analyze the relationship between FENet features and movement directions. PLSR projects the predictive and observed variables into a new space by finding pairs of weight vectors that maximize the covariance between the two projections [47]. Specifically, for each pair of neural activity data ($N$) and corresponding movement directions ($M$), we computed the FENet features from the neural activity in 30 Hz frequency bins and transformed these values into z-scores. After determining the weight matrix from the training dataset using PLSR, we used it to predict $M$ from $N$ on untrained data. Finally, we applied exponential smoothing to the predicted $M$ values.
@@
[…]
@@
Participants performed two reaction time tasks of increasing complexity (see Figures 3A and 3B). For both tasks, participants were required to react when they saw a visual stimulus. The tetraplegic participant BCI-JJ attempted a click movement as if pushing a button with one of his hand effectors via the BCI system, whereas each motor intact participant performed a click with their right index finger using a computer mouse. In both tasks, the timing of the stimulus was corroborated using a photodiode and the click signal was recorded directly from the voltage of the mouse button switch for the motor intact individuals. The participants sat 150 cm away from a computer monitor with a diagonal length of 54 cm.
The simple reaction time task required the participants to click as soon as they saw a white circular target with 5 cm radius appear on a black screen during the Target phase (see Figure 3A). The time of the appearance of the target was randomly taken from a uniform distribution between 1000 and 3000 ms. The target appeared on the screen for 60 ms. Each run of this task had 50 trials, with 10 randomly interleaved NO-GO catch trials during which the participants were not supposed to click, indicated by an auditory stimulus during the ITI phase.
The realistic braking task involved clicking to activate the brakes of a vehicle whenever an obstacle appeared in front of a vehicle in a virtual environment within the CARLA 0.9.13 driving simulator (see Figure 3B). During this task, the vehicle operated under algorithmic directional control at a constant velocity (5 mph). As the vehicle proceeded down a straight road, an obstacle (e.g., a vending machine, a trash can, etc.) would appear in the vehicle’s path (20 m in front of the vehicle) for each GO phase of the 40 trials per run. The participants were instructed to apply a brake signal using the BCI as soon as possible after seeing an obstacle appear. With no brake application, the vehicle would collide with the object in 5 sec. Coming to a full stop required 2 sec of constant brake application. Thus, the participant had 3 sec in which to begin applying the brake signal to avoid collision. A trial was considered successful if there was no collision during the GO phase and no manual braking during the NO-GO phase.
@@
The overlay was a Python pygame instance, which subscribed to the effector controller and transformed the x-value into the horizontal position of the blue circle and the y-value into the vertical position of the red circle. The overlay further post-processed the horizontal and vertical positions into steering and speed, respectively (for all three driving tasks) and the click signals into full-stop braking (for the last two bimanually controlled driving tasks). The steering value would change counterclockwise or clockwise exponentially when the blue circle moved to the left or right red rectangular hot zone. The speed value would increase (up to the maximum speed limit) or decrease (down to zero) exponentially when the red circle moved to the top or bottom red rectangular hot zone. If a full-stop braking was triggered by a click signal, the speed value would be immediately brought down to zero for 1 second, meaning the red circle would be immediately brought down to the bottom hot zone. For the two teledriving tasks without traffic, we also programmed slight changes of the steering value when the blue circle moved within the gray central cold zone. The final transformed steering and speed (and full-stop braking) values were displayed on screen and transmitted to the data client. The data client also sent the vehicle feedback values to the overlay display during the obstacle-course teledriving task.
@@
For the simulated driving task modified from the CARLA Autonomous Driving Leaderboard 2.0, we asked BCI-JJ and the group of 20 motor intact participants to hit the vehicle’s brake by clicking with their left index finger and to adjust the vehicle’s speed and steering by moving the cursor with their right thumb. While BCI-JJ had his attempted movement directly decoded from his PPC and MC signals using our BCI system, the group of 20 motor intact participants physically used a gaming joystick to control the agent vehicle’s movement (see figure S5). With the overlay instance floating above the vehicle simulation window (see Figure 3E), each participant got real-time visual feedback of their cursor and click control, the view ahead of the vehicle, and the side views. In the CARLA Leaderboard 2.0, we tested in Town 12’s downtown areas which featured high-rise skyscrapers arranged into blocks on a consistent grid of roads with busy traffic. The route that we edited from Town 12 included 2 right turns, 2 left runs, 1 right curve, 1 left lane switch, and 6 traffic lights (see Figure 3F). The correct route was labeled with green dots on the ground for the agent vehicle to follow.
The display computer recorded videos for all 10 runs in four sessions controlled by BCI-JJ and for both two runs in two sessions controlled by each motor intact participant (see movie S5 for a side-by-side comparison of the average simulated driving performance between BCI-JJ and a motor intact participant). Each run of the same route lasted for about 20 minutes with both the agent vehicle’s maximum speed and the traffic speed at 5 mph. If the agent vehicle deviated from the correct route for more than 10 seconds, the run would be stopped immediately and marked as incomplete. With the bimanual cursor-and-click control, the simulated driving performance through this route in the virtual town was evaluated based on (1) the proportion of the completed route distance ($C$), (2) the number of collisions with obstacles ($N_{c}$), (3) the number of lane deviations ($N_{l}$), and (4) the number of running red traffic lights ($N_{s}$). All these factors were automatically measured by the CARLA Leaderboard 2.0 by the end of each run and aggregated together for a composite bimanual simulated driving score as shown in Eq. 1. It started with an ideal 1.0 base score in each run, which was reduced each time an infraction was committed.
@@
As BCI-JJ was living in California while the Ford research team and this Mach-E vehicle were located in Michigan, it was necessary to establish a teleoperation framework in the initial stages of testing and validation. Our experimental setup consisted of a network created between the display computer and a modem in the Ford Mustang Mach-E SUV (see Figure 2B). A TailScale Virtual Private Network was set up to facilitate the generation of a different IP address for each of the different nodes. The display computer was connected to the Internet via a 1 Gbps fiber connection, while the modem on the vehicle was a Verizon MiFi-2100 5G Modem. In addition, a smartphone was securely mounted on the dash inside the vehicle to allow safety drivers to communicate with the team via video chat. This smartphone was not connected to the other devices in the vehicle to minimize network traffic. The video feed was transported to the display computer over WebRTC with a latency under 600ms and a bandwidth under 3Mbps. Additionally, the in-vehicle telematics node transmitted the actual speed and steering values to the display computer, which was then displayed in a rolling time window.
In any situation that requires real-time feedback for teleoperation, it is extremely important that command and video latencies are minimized. Therefore, we implemented a system for monitoring the glass-to-glass latency that involved initially pointing a camera at a clock in the vehicle, then streaming the feed to a display computer, and finally taking a picture of both the real time on the clock and the time as shown in the camera stream. The difference between these two times was the glass-to-glass latency. We ran the “date-time” Linux command in a terminal as part of the stream from the interface computer to the display computer. The display computer streamed its screen back to the vehicle via the external video chat.
@@
For the Mcity teledriving task with cursor-movement-only BCI control, we tested four randomly mapped routes in this mock urban environment. BCI-JJ navigated the Mach-E vehicle for one run per route across two sessions and finished each route between 2.5 minutes and 7.5 minutes. We used a camera mounted outside the front windshield of the vehicle to record full videos of those 4 runs (see movies S1S2S3 and S4), which were sent to three independent evaluators after each session to evaluate according to the infraction rubrics. The BCI teledriving infraction components in Mcity included (1) the proportion of the completed route distance ($C$), (2) the number of collisions with obstacles ($N_{c}$), and (3) the number of lane deviations ($N_{l}$). All these factors were aggregated together for a composite BCI teledriving score in Mcity as shown in Eq. 1. It started with an ideal 1.0 base score in each run, which was reduced each time an infraction was committed. $N_{s}$ in Eq. 1 was set to zero for this task, because no red traffic light or stop sign was considered.
@@
For the obstacle-course teledriving task with the cursor-and-click-based bimanual BCI control, we tested four loop-shaped route options containing many components of a standard driving test. Each route was clockwise or counterclockwise with or without a lane-switch segment. BCI-JJ navigated the Mach-E vehicle for 9 runs per route across three sessions and usually finished each run on this obstacle course between 2.5 minutes and 3 minutes. We used the camera mounted within the vehicle to record full videos for all 36 runs (see movies S6S7S8 and S9 for one example run per route), which were sent to three independent evaluators after each session to evaluate according to the infraction rubrics. The BCI teledriving infraction components through this obstacle course included (1) the proportion of the completed route distance ($C$), (2) the number of collisions with obstacles ($N_{c}$), (3) the number of lane deviations ($N_{l}$), and (4) the number of running stop signs ($N_{s}$). All these factors were aggregated together for a composite BCI teledriving score through the obstacle course as shown in Eq. 1. It started with an ideal 1.0 base score in each run, which was reduced each time an infraction was committed.

### v2

@@ DISCUSSION
The neural recordings in this study were collected from one participant enrolled in a BCI clinical study (ClinicalTrials.gov Identifier: NCT01958086). The institutional review boards of the California Institute of Technology, Casa Colina Hospital and Centers for Healthcare, and the University of California, Los Angeles, approved study procedures, including informed consent, implant surgery, and experiment design. BCI-JJ is a right-handed male 50 years old with a C4-C5 level spinal cord injury that occurred approximately three years before his enrollment in the study. BCI-JJ was implanted with two arrays of 96-channel NeuroPort Electrodes, including one array in the left superior parietal lobule of his PPC (MNI -34, -58, 67) and the other in his left MC close to the hand knob (MNI -39, -15, 69). Neural signals were acquired, amplified, bandpass filtered (0.3 Hz to 7.5 kHz), and digitized (30 kHz, 16 bits/sample) from electrodes using the neural signal processors called NeuroPort Systems (Blackrock Neurotech, Salt Lake City, UT).
It is important to note that BCI-JJ’s electrodes had been implanted for more than five years at the time of testing. The signal-to-noise ratio (SNR) of the recorded brain signals had decreased to approximately one-third of the value observed in the first year after implantation. Due to this decline, a custom feature extraction algorithm, FENet (see the “Feature extraction” section), played a crucial role in enabling BCI-JJ to control the vehicle with the proficiency described in this paper.
@@
We recruited 20 motor intact individuals, including 10 males and 10 females with an average age of $50\pm 15$, the same as the age of BCI-JJ at the time of testing. The motor intact control group participated in the simple reaction time task, the braking reaction time task, and the simulated town driving task. Each motor intact participant attended two sessions on their own to complete these three tasks. All participants were informed about the purpose of the study and provided their written informed consent prior to participating. They were compensated at the end of each session based on their task performance. 11 out of 20 motor intact participants filled out an online post-session questionnaire after they completed the simulated town driving task. All those 11 people are right-handed. They all got their regular driver’s licenses before participating. The rest of the questionnaire’s results are shown in figure S8. Although $36.4\%$ of them would experience motion sickness when they take a ride or even drive a physical vehicle (see figure 8A), none of them experienced motion sickness when they controlled the virtual vehicle in our simulated setting (see figure 8B). Regarding their past experiences in simulated driving or other simulated games, $63.6\%$ of them never tried simulated driving, but $54.5\%$ had some experiences in other simulated games (see figure 8C). $90.9\%$ of them had gaming joystick control experience(s) at least once or twice before participating in this task (see figure 8D) and $54.6\%$ of them found the gaming joystick control in this task relatively easy (see figure 8E). $72.7\%$ of them did not feel fatigued during this task, and $18.2\%$ only felt occasionally or towards the end, with the rest $9.1\%$ could not recall their fatigue level (see figure 8F).
@@
We used PLSR to analyze the relationship between FENet features and movement directions. PLSR projects the predictive and observed variables into a new space by finding pairs of weight vectors that maximize the covariance between the two projections [51]. Specifically, for each pair of neural activity data ($N$) and corresponding movement directions ($M$), we computed the FENet features from the neural activity in every 30 ms and transformed these values into z-scores. After determining the weight matrix from the training dataset using PLSR, we used it to predict $M$ from $N$ on untrained data. Finally, we applied exponential smoothing to the predicted $M$ values.
@@
Cursor-and-click training
BCI-JJ was asked to perform a series of center-out task runs to train our cursor-movement decoder and our click classifier in each session. The center-out task involved a central fixation point and eight radial targets spaced at 45-degree intervals around a circle. Each run consisted of 48 trials. In each trial, one of the targets would be randomly selected to change its color from gray to red, signaling the start of the trial. A cursor would then move from the center to the indicated target. After the cursor reached the target, it would return to the center.
BCI-JJ completed three runs of the center-out task to train the cursor-movement decoder for each driving session. In this case, the participant was instructed to attempt to move his right thumb following the cursor movement. After each run, the decoder was improved using data from all previous runs. Each subsequent center-out task had decreasing assistance levels: 1.0, 0.4, and 0.2. An assistance level of 1.0 indicated that the computer was fully controlling the cursor with no input from the participant, while an assistance level of 0 meant the participant had complete control over the cursor in a closed-loop test run. During the training and testing process, there sometimes would appear a large decoder bias of cursor movement pulling in the horizontal or vertical direction, which was different from human control error. BCI-JJ would let us know when it happened so that we could adjust the anti-bias setting of the decoder accordingly.
To evaluate the decoder performance, we used the cross-validated coefficient of determination ($R^{2}$) as a measure of the strength of the linear association between the predicted and intended kinematics:
$R^{2}=\frac{\sum_{i=1}^{n}(y_{i}-\hat{y_{i}})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}.$ (2)
Here $n$ was the number of observations. For each observation $i$, $y_{i}$ was the observed kinematics, $\hat{y_{i}}$ was the predicted kinematics. $\bar{y}$ was the mean of all observed kinematics. The $R^{2}$ values for the horizontal and vertical dimensions were computed independently and then averaged for reporting.
BCI-JJ performed five runs of a modified center-out click task in the open loop to train the click classifier for each reaction-time or simulated town driving session. In this task, the cursor would briefly change color when reaching a target on $70\%$ of trials. BCI-JJ was asked to attempt clicking, such as flexing his left index finger, when the cursor changed color. For each reaction-time session that only relied on clicks, the assistance level of the center-out click task remained at 1.0 for computer-controlled cursor movement. For each simulated driving session that relied on bimanual cursor-and-click control, the click classifier was trained after completing the cursor-movement decoder training, with the cursor assistance level set to 0.2. The trained click classifier was tested during a subsequent closed-loop run of the same center-out click task, in which an attempted click would cause the cursor to change color as feedback, while BCI-JJ maintained complete control over cursor movement if the cursor-movement decoder was trained earlier in the same session.
Lag estimation
For each driving task, we estimated the system’s lag from the recorded videos of each session by calculating the maximum lag between the participant’s intended movement and the actual visual feedback of the vehicle moving in the intended direction (see figure 9A). We identified the participant’s intention by calculating the steering cursor’s position on the overlay for each video frame, by locating blobs of blue pixels and averaging their positions. We estimated the vehicle movement by calculating optical flow from the video using the RAFT deep learning algorithm [52]. This method t+akes two images as input and computes the displacement of each pixel in the first image to its corresponding pixel in the second, generating an array of vectors representing the movement of each pixel. The analyzed video had a resolution of $1920\times 980$ and was converted to grayscale for the analysis. The optic flow method was applied to a smaller area of $800\times 141$ pixels from the original video, which had an unobstructed view of the road and horizon. The vehicle movement velocity vector was calculated across frames every 66 ms as the mean of the magnitudes of all individual optic flow vectors whose magnitudes exceeded a predefined threshold. We quantified the difference between video frames by calculating the structural similarity index measure (SSIM) between two consecutive frames. Both the participant’s intention and the video feedback of the vehicle movement were smoothed to remove noise. Finally, we calculated the cross-correlation between the participant’s intention and the video feedback of the vehicle movement and identified the maximum lag (see figures 9B and 9C as an example from the Mcity teledriving task). The cross-correlation was calculated for frames with an SSIM less than 0.99.
We calculated an adjusted score due to the varying lengths of the runs. This adjusted score was determined using the following equation:
$\text{score}_{adj}=\text{score}_{org}\times\text{length\_rate}.$ (3)
Here scoreorg was the original score of a run, and length_rate was computed as the route length of the current run divided by the mean of the route lengths of all runs (see table S12 for a summary of route lengths for three driving tasks). Then we could calculate the Pearson correlation between the magnitude of the lags and the adjusted scores.
This method of measuring lags accounted for all lags between the participant’s decoded intention and the visual feedback in our system during execution of that motor plan, except for those in the decoder process and in the transmission of that information to the display PC. The uncomputed lags were estimated to be negligible compared to the previously analyzed lag, because the decoder processed the recordings in 30-ms bins, and the digital communication delay between the decoder and the display PCs was calculated to be less than 1 ms using the ping utility to measure LAN latency.
Channel contributions to decoder
We calculated the Shapley values for each channel to determine their contributions to the decoding process. The Shapley value of a feature is determined by the deviation in the decoder’s output at a specific query point from its average output when the feature is excluded [53]. Features with higher Shapley values are considered more important because they have a greater impact on the decoder’s output.
To calculate the Shapley values, we first computed the mean of the FENet features for each channel and time point, and trained our decoder using a procedure similar to that described in the “Cursor-and-click training” section. Afterwards, we determined the mean of the Shapley values across 120 randomly selected query points for each channel. The contribution of each brain region was calculated as the average of all Shapley values associated with the channels corresponding to that brain area, divided by the total sum of the Shapley values for all channels.
@@
We also performed a power analysis to establish the required sample size of individuals with intact motor function for comparisons between two independent means. The analysis was based on an effect size of 0.8, an alpha level of 0.05, a power of 0.95, considering a dropout rate of $25\%$.
@@
Participants performed two reaction time tasks of increasing complexity (see Figures 3A and 3B). For both tasks, participants were required to react when they saw a visual stimulus. The tetraplegic participant BCI-JJ attempted a click movement as if pushing a button with one of his hand effectors via the BCI system, whereas each motor intact participant performed a click with their right index finger using a computer mouse. In both tasks, the timing of the stimulus was corroborated using a photodiode, and the click signal was recorded directly from the voltage of the mouse button switch for the motor intact individuals. The participants sat 150 cm away from a computer monitor with a diagonal length of 54 cm.
The simple reaction time task required the participants to click as soon as they saw a white circular target with a 5-cm radius appear on a black screen during the Target phase (see Figure 3A). The onset of the target was randomly taken from a uniform distribution between 1000 and 3000 ms from the start of a trial. The target appeared on the screen for 60 ms. A reaction time comparison to the previous trial would appear on the screen after 1000 ms from the target onset for the current trial; therefore, a valid click must have a reaction time within 1000 ms. However, a click within the first 50 ms from the target onset would be counted as a too-soon false positive click, because such a click might reflect a target prediction from the participant according to the pace of the trials instead of an actual reaction to the new target. Each run of this task had 50 trials, with 10 randomly interleaved NO-GO catch trials during which the participants were not supposed to click, indicated by a low-pitch auditory stimulus during the ITI phase. The valid reaction time was measured between 50 ms and 1000 ms for GO trials.
The braking reaction time task involved clicking to activate the brakes of a vehicle whenever an obstacle appeared in front of a vehicle in a virtual environment within the CARLA 0.9.13 driving simulator (see Figure 3B). During this task, the vehicle operated under algorithmic directional control at a constant velocity (5 mph). As the vehicle proceeded down a straight road, an obstacle (e.g., a vending machine, a trash can, etc.) would appear in the vehicle’s path (20 m in front of the vehicle) for each GO phase of the 40 trials per run. The participants were instructed to brake using the BCI-decoded click signal as soon as possible after seeing an obstacle appear. Without brake application in time, the vehicle would collide with the obstacle in front. Coming to a full stop required a one-time click that is programmed for a 2-second constant brake. The participant had 1 second from the obstacle onset to begin braking to avoid a collision; therefore, a valid reaction time must be within 1000 ms. A trial was considered successful if there was no collision during the GO phase and no manual braking during the NO-GO phase. However, a click within the first 50 ms from the obstacle onset would be counted as a too-soon false positive click even if there was no collision, because such a click might reflect an obstacle prediction from the participant according to the pace of the trials instead of an actual reaction to the new obstacle. The valid reaction time was measured between 50 ms and 1000 ms for GO phases.
@@
The overlay was a Python pygame instance, which subscribed to the effector controller and transformed the x-value into the horizontal position of the blue circle and the y-value into the vertical position of the red circle. The overlay further post-processed the horizontal and vertical positions into steering and speed, respectively (for all three driving tasks) and the click signals into full-stop braking (for the bimanually controlled driving task in the simulator). The steering value would change counterclockwise or clockwise exponentially when the blue circle moved to the left or right red rectangular hot zone. The speed value would increase (up to the maximum speed limit) or decrease (down to zero) exponentially when the red circle moved to the top or bottom red rectangular hot zone. If a full-stop braking was triggered by a click signal, the speed value would be immediately brought down to zero for 1 second, meaning the red circle would be immediately brought down to the bottom hot zone. For the two teledriving tasks without traffic, we also programmed slight changes of the steering value when the blue circle moved within the gray central cold zone. The final transformed steering and speed (and full-stop braking) values were displayed on screen and transmitted to the data client. The data client also sent the vehicle feedback values to the overlay display during the obstacle-course teledriving task.
@@
For the simulated town driving task modified from the CARLA Autonomous Driving Leaderboard 2.0, we asked BCI-JJ and the group of 20 motor intact participants to hit the vehicle’s brake by clicking with their left index finger and to adjust the vehicle’s speed and steering by moving the cursor with their right thumb. While BCI-JJ had his attempted movement directly decoded from his PPC and MC signals using our BCI system, the group of 20 motor intact participants physically used a gaming joystick to control the agent vehicle’s movement (see figure S6). With the overlay instance floating above the vehicle simulation window (see Figure 3G), each participant got real-time visual feedback of their cursor and click control, the view ahead of the vehicle, and the side views. In the CARLA Leaderboard 2.0, we tested in Town 12’s downtown areas which featured high-rise skyscrapers arranged into blocks on a consistent grid of roads with heavy traffic. The route that we edited from Town 12 included 2 right turns, 2 left runs, 1 right curve, 1 left lane switch, and 6 traffic lights (see Figure 3H). The correct route was labeled with green dots on the ground for the agent vehicle to follow.
The display computer recorded videos for all 10 runs in four sessions controlled by BCI-JJ and for both two runs in two sessions controlled by each motor intact participant (see movie S9 for a side-by-side comparison of the average simulated town driving performance between BCI-JJ and a motor intact participant). Each run of the same route lasted for about 20 minutes with both the agent vehicle’s maximum speed and the traffic flow speed at 5 mph. If the agent vehicle deviated from the correct route for more than 10 seconds, the run would be stopped immediately and marked as incomplete. With the bimanual cursor-and-click control, the simulated driving performance through this route in the virtual town was evaluated based on (1) the proportion of the completed route distance ($C$), (2) the number of collisions with obstacles ($N_{c}$), (3) the number of lane deviations ($N_{l}$), and (4) the number of running red traffic lights ($N_{s}$). All these factors were automatically measured by the CARLA Leaderboard 2.0 by the end of each run and aggregated together for a composite bimanual simulated town driving score as shown in Eq. 1. It started with an ideal 1.0 base score in each run, which was reduced each time an infraction was committed.
@@
As BCI-JJ was living in California while the Ford research team and this Mach-E vehicle were located in Michigan, it was necessary to establish a teleoperation framework in the initial stages of testing and validation. Our experimental setup consisted of a network created between the display computer and a modem in the Ford Mustang Mach-E SUV (see Figure 2B). A TailScale Virtual Private Network was set up to facilitate the generation of a different IP address for each of the different nodes. The display computer was connected to the Internet via a 1 Gbps fiber connection, while the modem on the vehicle was a Verizon MiFi-2100 5G Modem. In addition, a smartphone was securely mounted on the dash inside the vehicle to allow safety drivers to communicate with the team via video chat. This smartphone was not connected to the other devices in the vehicle to minimize network traffic.
@@
For the Mcity teledriving task with single-effector BCI control over the cursor movement, we tested four randomly mapped routes in this mock urban environment without following any traffic lights or stop signs. BCI-JJ navigated the Mach-E vehicle for one run per route across two sessions and finished each route between 2.5 minutes and 7.5 minutes. We used a camera mounted outside the front windshield of the vehicle to record full videos of those 4 runs (see movies S1S2S3 and S4), which were sent to three independent human evaluators after each session to evaluate according to the infraction rubrics. The BCI teledriving infraction components in Mcity included (1) the proportion of the completed route distance ($C$), (2) the number of collisions with obstacles ($N_{c}$), and (3) the number of lane deviations ($N_{l}$). All these factors were aggregated together for a composite BCI teledriving score in Mcity as shown in Eq. 1. It started with an ideal 1.0 base score in each run, which was reduced each time an infraction was committed. $N_{s}$ in Eq. 1 was set to zero for this task, because no red traffic light or stop sign was considered.
@@
For the obstacle-course teledriving task with the single-effector BCI control over the cursor movement, we tested four loop-shaped route options containing many components of a standard driving test. Each route was clockwise or counterclockwise with or without a lane-switch segment. BCI-JJ navigated the Mach-E vehicle for 9 runs per route across three sessions and usually finished each run on this obstacle course between 2.5 minutes and 3 minutes. We used the camera mounted within the vehicle to record full videos for all 36 runs (see movies S5S6S7 and S8 for one example run per route), which were sent to three independent evaluators after each session to evaluate according to the infraction rubrics. The BCI teledriving infraction components through this obstacle course included (1) the proportion of the completed route distance ($C$), (2) the number of collisions with obstacles ($N_{c}$), (3) the number of lane deviations ($N_{l}$), and (4) the number of running stop signs ($N_{s}$). All these factors were aggregated together for a composite BCI teledriving score through the obstacle course as shown in Eq. 1. It started with an ideal 1.0 base score in each run, which was reduced each time an infraction was committed.
Safety criteria for driving tasks
We set up the safety criteria for all three driving tasks to include basic components of a standard driving test for motor intact individuals, in accordance with part of the evaluation and metrics from CARLA Autonomous Driving Leaderboard 2.0 and part of the checklist for stops, lane use, lane change, and vehicle control in the “Driving Performance Evaluation Score Sheet” [44] from the California Department of Motor Vehicles (DMV) (see table S13). Different from the California DMV score sheet, we did not apply any turn signal light before lane changes or turns in any of the three driving tasks, and we did not consider yielding in the traffic scenario in the simulated town driving task. Different from the Leaderboard 2.0 evaluation and metrics, we did not add any emergency vehicles to the scene, so we did not measure failure to yield to an emergency vehicle in the simulated town driving task. Both the commercially available Ford Mustang Mach-E vehicle and the virtual agent vehicle were configured to have a slow maximum speed (i.e., 4 mph and 5 mph, respectively), which was the same as the traffic flow speed in the simulated case. When a participant controlled the cursor’s vertical movement on the overlay via either BCI-decoded signals or the gaming joystick, the vehicle speed could always be kept within a safe range. We also did not introduce pedestrians in any of the three driving tasks for easier safety regulations. With steering and speed control for the teledriving tasks plus additional full-stop braking control for the simulated town driving task, both BCI-JJ and the motor intact control group were supposed to complete each driving run with fewer occurrences of the infractions (including collisions, lane deviations and running red traffic lights or stop signs) to obtain safer driving performance as evaluated by the three human evaluators for the teledriving tasks and by the CARLA Leaderboard 2.0 itself for the simulated town driving task.